OpenStudy (anonymous):

use elimination: please help and show steps so i can understand! 2.5x+2y=10.7 -2.5x+2y=12.9

OpenStudy (superdavesuper):

Elimination means to eliminate a variable in order to solve a prob. In this case, add the two equations together and see what happens plz

OpenStudy (anonymous):

yeah nevermind i figured this one out. lol sorry.

OpenStudy (anonymous):

well wait actually, i know you have to subtract the y but idk what next

OpenStudy (superdavesuper):

OK after subtraction wats left on the left n right of the = sign?

OpenStudy (anonymous):

and i wrote the problem wrong at first this is the right one: 2.5x+y=10.7 -2.5x+2y=12.9

OpenStudy (superdavesuper):

Actually a simpler way is just add them together n see wats left. try it :)

OpenStudy (anonymous):

i don't know… 3y=23.6?

OpenStudy (superdavesuper):

Yup so y equals to.....?

OpenStudy (anonymous):

7.8?

OpenStudy (superdavesuper):

i got 7.9 but close enough lol now solve for x n u are DONE :)

OpenStudy (anonymous):

well I'm not sure how to do that.

OpenStudy (superdavesuper):

put y=7.9 back into either 1 of da eqn

OpenStudy (anonymous):

ok so, 2.5x+7.9=10.7 -7.9 -7.9 2.5x=2.8 _________ 2.5 2.5 x=1.12 correct?

OpenStudy (agent0smith):

^yep.